3. A market demand is given by QA = 200 – 3PA + 2PB + 0.04M, where QA is the quantity demanded of good A, PA is good A’s price, PB is good B’s price, and M is the average good A’s consumer monthly income. If PA = $12, PB = $10, and M = $2,500,

a. calculate good A’s (point) own-price elasticity;

b. calculate good A’s (point) cross-price elasticity;

c. calculate good A’s (point) income elasticity;
1
Expert's answer
2019-05-27T11:59:16-0400

If PA = $12, PB = $10, and M = $2,500, then:

QA = 200 – 3PA + 2PB + 0.04M = 200 – 36 + 20 + 0.04×2,500 = 284 units.

The formula for point-price elasticity is:

Ed = b×P/Q.

a. good A’s (point) own-price elasticity is:

Ed = -3×12/284 = - 0.13.

b. good A’s (point) cross-price elasticity is:

Ecp = 2×10/284 = 0.07.

c. good A’s (point) income elasticity is:

Ei = 0.04×2500/284 = 0.35.